Suppose the U.S. president wants an estimate of the proportion of the population who support his current policy toward revisions in the Social Security system. The president wants the estimate to be within 0.026 of the true proportion. Assume a 98 percent level of confidence. The president's political advisors estimated the proportion supporting the current policy to be 0.61. (Round up your answers to the next whole number.) (a) How large of a sample is required? . (b) How large of a sample would be necessary if no estimate were available for the proportion that support current policy? .
Answer #1

a)

Here, p = 0.61


ME = 0.026

z value at 98% confidence interval is,

alpha = 1 - 0.98 = 0.02
alpha/2 = 0.02/2 = 0.01
Zalpha/2 = Z0.01 = 2.3263

By using Me formula, we get

ME = z *sqrt(p*(1-p)/n)

0.026 = 2.3263 *sqrt(0.61 *(1-0.61)/n)
n = (2.3263/0.026)^2 *0.61*(1-0.61)
n = 1904.5705

Rounding to nearest whole no
n = 1905

b)

a)

Here, p is not given by assuming p = 0.5

ME = 0.026

z value at 98% confidence interval is,

alpha = 1 - 0.98 = 0.02
alpha/2 = 0.02/2 = 0.01
Zalpha/2 = Z0.01 = 2.3263

By using Me formula, we get

ME = z *sqrt(p*(1-p)/n)

0.026 = 2.3263 *sqrt(0.5 *(1-0.5)/n)
n = (2.3263/0.026)^2 *0.5*(1-0.5)
n = 2001.440
Rounding to nearest whole no
n = 2001